can someone solve this problem for me?
find two consecutive even integers such that smaller added to two times the larger gives a sum of 34

Answer by richwmiller(17219)

x+2*(x+2)=34
Solve for x
"x" is the smaller even integer.
"x+2" is the larger
x+2*(x+2)=34
x+2x+4=34
You can finish it, can't you?
Incredible! The student says "This is no help at all!"